OpenStudy (anonymous):

fan and medal for best answer :) an expression is shown below: f(x)=-16x^2+8x+3 part a.) what are the x intercepts of the graph of f(x)? show work part b.) is the vertex of the graph of f(x) going to be maximum or minimum? what are the coordinates of the vertex? justify answer and show work part c.) what are the steps you would use to graph f(x)? make sure you can use answer obtained in part a and part b to draw graph.

OpenStudy (hoblos):

x intercepts are values of x for y=0 so -16x^2+8x+3 = 0 can you solve this quadratic equation ?

OpenStudy (hoblos):

we will solve them one by one so let's start with the first do you know the quadratic formula ?

OpenStudy (31356):

@hoblos is correct! :D

OpenStudy (31356):

Yes, that's the quadratic formula. Go to this (for more help) http://www.purplemath.com/modules/quadform.htm

OpenStudy (31356):

So first, do you know what x-intercept is?

OpenStudy (hoblos):

@31356 thanks :) @bangbang559 the quadratic formula to solve this quadratic equation is \[x= \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (31356):

My pleasure :D Thanks for the medal too :D

OpenStudy (31356):

You deserve a medal too! :D

OpenStudy (31356):

Go to the website, it has some given examples too. :D

OpenStudy (31356):

It does seem confusing at the first place, however practice makes perfect, lol :D

OpenStudy (31356):

If you don't practice then you're not able to know how to solve problems similar to this :D

OpenStudy (31356):

Here ask @satellite73 for help. :D

OpenStudy (31356):

He's good! :D

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

@secret66

OpenStudy (anonymous):

@SnuggieLad

OpenStudy (anonymous):

@AuroraB

OpenStudy (anonymous):

@MathLegend

OpenStudy (anonymous):

Refer to the Mathematica calculation attached.

OpenStudy (anonymous):

@robtobey witch answer is this?

OpenStudy (anonymous):

@bangbang559 Have you viewed the attachment, bangbang559_2.pdf ? All of the answers are there.

OpenStudy (anonymous):

where are the x intercepts of the graph ? @robtobey

OpenStudy (anonymous):

x=-1/4 and x=3/4

OpenStudy (anonymous):

Thank you for the medal and have a good day.

OpenStudy (anonymous):

anytime :) and you too